OpenStudy (anonymous):

cot(x-pi/2) -tanx is undercofunction identities, but it's positive how do i turn it negative

OpenStudy (anonymous):

@precal

OpenStudy (precal):

no clue, don't know anything about cofunction off the top of my head.....

OpenStudy (precal):

would have to go look it up and I misplaced my precal book, trying to clean my room

OpenStudy (anonymous):

aww okay thanks though

OpenStudy (anonymous):

that is what a cofunction identity is

OpenStudy (precal):

listen to satellite73

OpenStudy (anonymous):

oh i see the confusion

OpenStudy (anonymous):

but my question is to verify cot(x-pi/2)=-tanx

OpenStudy (anonymous):

\[\cot(\frac{\pi}{2}-x)=\tan(x)\]

OpenStudy (anonymous):

that is a cofunction identity

OpenStudy (anonymous):

the identity only has cot(pi/2-x)=tanx

OpenStudy (anonymous):

right

OpenStudy (anonymous):

hint: cotangent is odd

OpenStudy (anonymous):

okay ill try it

OpenStudy (anonymous):

ok but there is nothing much to try it comes from the fact that cotangent is an odd function and \(x-\frac{\pi}{2}=-(\frac{\pi}{2}-x)\)

OpenStudy (anonymous):

so i just make it negative? @satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\cot(-\theta)=-\cot(\theta)\]

OpenStudy (anonymous):

@satellite73 it doesn't make sense to me

OpenStudy (precal):

those are called even and odd identities